OpenStudy (anonymous):

Hola! This wording is so confusing... I know how to solve this type of problem but "the difference" makes me wonder if they want subtraction or a negative variable. Thanks for any help! I DO NOT want the answer to the set-up, just how to get my equation set up in the first place :) The difference of a number times 4 and 3 equals 5

OpenStudy (asnaseer):

You are told: "The difference of a number times 4 and 3 equals 5" break this up into: "The difference of" "a number times 4" "and" "3" "equals" "5" does that help?

OpenStudy (anonymous):

n4-n3=5 ?

OpenStudy (asnaseer):

If I said: "the difference of 18 and 8 is 10" it means: 18 - 8 = 10

OpenStudy (anonymous):

n-(4+3)=5 ? I think your getting me on the right track!

OpenStudy (asnaseer):

not quite :)

OpenStudy (asnaseer):

lets break this down further, we stat with: "The difference of" "a number times 4" "and" "3" "equals" "5" lets call the unknown number "n" then we can replace "a number times 4" with 4n agreed so far?

OpenStudy (anonymous):

okay, that makes sense...

OpenStudy (asnaseer):

so we now have: "The difference of" "4n" "and" "3" "equals" "5" can you do it now?

OpenStudy (anonymous):

I feel so stupid... so its 4n-3=5?

OpenStudy (asnaseer):

perfect! and you are certainly not stupid my fried - the question is not very well worded :)

OpenStudy (anonymous):

You are awesome!!! Thanks for the encouragement--I guess not freaking out and breaking it down is helpful.

OpenStudy (anonymous):

something to apply to life, haha

OpenStudy (asnaseer):

exactly - keep up the positivity! :)